OpenStudy (anonymous):

The area of a trapzoid is 11 square meters one base is 3 meters long and the other is 8 meters long. Find the height of the trapazoid. Without substitution, solve the formula for the unknown variable in terms of the known variables.

OpenStudy (anonymous):

\[A=\frac{(b _{1}+b _{2})*h}{2}\]\[11=\frac{(3+8)*h}{2}\] 11*2=11*h h=22/11 h=2 If you like my answer click good answer,ty